OpenStudy (anonymous):

Use differentials to approximate the amount in cubic feet of paint needed to cover a cube of side length 100 feet if thickness of the paint is to be 0.01 feet.

OpenStudy (bahrom7893):

i dont remember.. any formulas to remind me?

OpenStudy (anonymous):

can you go to this link and look at question 9 http://www.math.sunysb.edu/~jwalsh/Summer_08/hmwk0625sol.pdf

OpenStudy (anonymous):

its in pdf so i cant copy and paste but question 9 is a sample problem

OpenStudy (bahrom7893):

ok

OpenStudy (bahrom7893):

easy, unless im very wrong: V = a^3 dV = 3a^2(da) = 3*100^2*(0.01) = 3*10,000*0.01 = 3*100 = 300

OpenStudy (anonymous):

ok thanks